# The rising costs of training frontier AI models

### v1

@@ A.1 Hardware acquisition cost
To estimate the acquisition cost of the hardware used for each training run, we used the following formula, where the intra-server cost per GPU includes the cost of CPUs, memory, intra-server networking (i.e. GPU-to-GPU), and server markup:
@@
To estimate accelerator costs, we matched the hardware for each training run with an estimated release price of the server. If the hardware type was a Google TPU, we simply looked up the accelerator cost calculated in Section A.2. Otherwise, we looked up the earliest known price corresponding to the accelerator. If the price was for a DGX server, we used that as the combined accelerator and intra-server hardware cost (dividing by the number of GPUs per server). [^2] HGX servers are more suited to large-scale, customized AI supercomputers, but we found very little information on their pricing, so we used DGX pricing. Otherwise, we found the earliest known price for an individual GPU and adjusted that price for additional server costs.
We calculated a server-to-GPU cost ratio based on known DGX and single-GPU prices near release, using the formula (DGX cost)/(8$\times$×GPU cost). We were able to estimate this for NVIDIA (1.54×)(1.54$\times$)( 1.54 × ), V100 (1.69×)(1.69$\times$)( 1.69 × ), and A100 (1.66×)(1.66$\times$)( 1.66 × ). For other NVIDIA chips, and for TPUs we used the mean of these three known factors (1.64×)(1.64$\times$)( 1.64 × ). We assumed that the DGX server prices (or equivalent) included the cost of intra-server interconnect switches and transceivers (e.g. NVLink). [^3] Some of the product pages where we found hardware prices listed NVSwitches as components, but it was unclear whether NVLink cables for intra-server links were included. We did not account for financing, i.e. the interest paid on a loan to purchase the hardware up-front.
To account for the cost of inter-server (i.e. server to server) networking, we used an estimate by Kostovic (forthcoming), based on the reference architecture of the NVIDIA H100 SuperPOD [15]. According to this estimate, approximately 19% of the total cost of the SuperPOD goes towards cluster-level interconnect for configurations with less than 4096 GPUs, and 20% for 4096 GPUs and above (due to the additional third layer of switches). An expert in AI hardware reported an estimated range of 10% to 20% for A100-based clusters with Infiniband [16].
Based on this, we used a constant overhead of 19% to account for cluster-level interconnect costs. We also assumed that this generalizes to TPU servers. However, we have considerable uncertainty in this. Different cluster architectures, and smaller or larger architectures, would have a different proportion of cost.
@@ A.3 Amortization model
Initial hardware value=Hardware acquisition $\text{Initial hardware value}=\dfrac{\text{Hardware acquisition cost}}{\exp\Big{(}\big{[}\text{Training start date}-\text{Hardware availability date}\big{]}\times\text{Depreciation factor}\Big{)}}$
The hardware availability date depended on the type of hardware. If the hardware was a Google TPU, we used the hardware announcement date. For GPUs, we used a 90-day buffer between the GPU first going on the market and the GPU actually being shipped to the buyer. Our results are robust to variations in this buffer time—see Section B.4.
@@
The precise way to amortize cost with exponential depreciation as follows:
[…]
$=\text{Initial hardware value}\times\text{Hardware quantity}\times\Big{(}1-\exp\big{[}-\text{Training time}\times\text{Depreciation factor}\big{]}\Big{)}$
However, we could estimate chip-hours more often and more reliably than the training time or hardware quantity separately. This is because chip-hours can also be estimated from training compute in FLOP divided by the FLOP/s achieved during training. We used a linear approximation to take advantage of these chip-hour estimates:
@@
Due to NVIDIA covering defects and component failures under warranty, we concluded that hardware failures are not a significant source of depreciation relative to hardware progress. As one data point, an average of 1 to 2 failures per week occurred when training the BLOOM model on a cluster of 384 NVIDIA A100 GPUs [25]. If these were all catastrophic failures, the expected hardware lifetime would be 3.7 years. We expect that NVIDIA replaces or repairs the GPUs on a much faster timescale, which makes the cost of failure small compared to hardware price depreciation.
@@ A.5 Cloud price selection
To estimate training costs from cloud rental prices, we matched the hardware type and publication date of each ML model with a price from the hardware price database. To do this, we first filtered the database for prices which matched the hardware type and the most likely cloud provider that would be used, with the latter based on the developer of the ML model, e.g. using Google Cloud for any Google lab, and using Microsoft Azure for OpenAI based on their partnership with Microsoft. We then estimated the date of hardware procurement as the publication date minus the training time, and minus a further 2 months to account for preparation before the training run. [^4] The choice of 2 months was an educated guess. The matching of models to cloud prices was not very sensitive to this choice because the price data was sparse and stable over time. If the training time was unavailable, we used the median duration of models with known training time, which was approximately 33 days.
@@ A.6 Accounting for compute used throughout model development
One source of evidence on the allocation of compute is the training of smaller model sizes for a given architecture. For example, smaller versions of GPT-3 used 4.5e22 FLOP (based on $compute=6\times parameters\times tokens$[13, Table 2.1]. This shows at least 14% of compute was spent outside the main training run. Similar reasoning for BLOOM reveals about 63% of compute was used on smaller models [25, Table 5].
@@ B.5 Excluding equity from R&D staff costs
To measure the impact of equity on the total amortized model development cost, Figure 9(a) shows the cost breakdown with equity excluded from the R&D staff cost. The proportion of cost on R&D staff decreases from about 30–50% with equity included, to 20–30% with equity excluded.
Figure 9: (top) Breakdown of total amortized model development costs for selected models, with equity excluded from the R&D staff cost. Hardware costs are amortized to the total number of chip-hours spent on experiments and training, while R&D staff costs cover the duration of development from initial experiments to publication. Error bars indicate 90% credible intervals, while the main bar values are medians. (bottom) Costs components as a percentage of the total, based on median estimates.
@@ Appendix C Power capacity for model training
where hardware peak TDP accounts for full server hardware. We find a growth rate of 2.0x per year (95% CI: 1.7x to 2.4x). Gemini Ultra has the largest estimated power capacity, at around 35 MW. Projecting this trend forward from Gemini Ultra, the most power-intensive training run would draw 1 GW by 2029. To put this in context, the top ten largest power plants in the United States have a capacity ranging from 3 GW to 7 GW [38].

### v2

@@ A.1 Hardware acquisition cost
The frontier AI models in our dataset were trained on clusters of many GPU or TPU chips (or “chips” for short). We set out to estimate the total cost of the chips, servers, and networking equipment in these clusters, which we call the hardware acquisition cost. This cost is calculated as follows:
@@
Where “Acquisition cost per chip” accounts for the GPU or TPU chip itself, other server costs (CPUs, memory, chip-to-chip networking, and markup), and the cost of server-to-server networking equipment. Table 2 shows how we calculated this quantity depending on what was known about the training hardware.
Known information Formula for “Acquisition cost per chip”
[…]
Table 2: The formula to estimate “Acquisition cost per chip”, depending on what is known about the training hardware.
If the training hardware was a GPU, we looked up the earliest known price linked to that GPU. If the price we found was for a single GPU, we multiplied that price by a “chip-to-server” cost factor (detailed later). If the price we found was instead for a DGX server, we divided that price by the number of GPUs per server. [^3] HGX servers are more suited to large-scale, customized AI supercomputers, but we found very little information on their pricing, so we used DGX pricing. Finally, if the training hardware was a Google TPU, we used the “Geometric mean” production cost estimate from Table 3. This represents the cost of the TPU chip itself, which we then multiplied by a chip-to-server cost factor.
We calculated chip-to-server cost factors based on known DGX and single-GPU prices near release, using the formula (DGX cost)/(8$\times$×GPU cost). We were able to estimate this for the NVIDIA P100 (1.54×)(1.54$\times$)( 1.54 × ), V100 (1.69×)(1.69$\times$)( 1.69 × ), and A100 (1.66×)(1.66$\times$)( 1.66 × ). For other NVIDIA chips, and for TPUs, we used the mean of these three known factors (1.64×)(1.64$\times$)( 1.64 × ). We assumed that the DGX server prices included the cost of chip-to-chip interconnect switches and transceivers. [^4] Some of the NVIDIA product pages where we found hardware prices listed NVSwitches as components, but it was unclear whether NVLink cables for chip-to-chip links were included. We did not account for financing, i.e. the interest paid on a loan to purchase the hardware up-front.
Once we had the cost per chip for a single server, we added the cost of server-to-server networking equipment. We used an estimate by Kostovic (forthcoming), based on the reference architecture of the NVIDIA H100 SuperPOD [15]. According to this estimate, approximately 19% of the total cost of the SuperPOD goes towards cluster-level interconnect for configurations with less than 4096 GPUs, and 20% for 4096 GPUs and above, due to an additional third layer of switches. Consistent with these figures, another expert in AI hardware estimated a range of 10% to 20% for A100-based clusters with an Infiniband network [16].
[…]
@@ A.3 Amortization model
Start value per chip=Acquisition cost per $\text{Start value per chip}=\dfrac{\text{Acquisition cost per chip}}{\exp\Big{(}\big{[}\text{Training start date}-\text{Hardware availability date}\big{]}\cdot r\ln{10}\Big{)}}$
where $r$ is a depreciation rate in orders of magnitude per year, and the difference in dates is in years. The hardware availability date depended on the type of hardware. If the hardware was a Google TPU, we used the hardware announcement date. For GPUs, we used a 90-day buffer between the GPU first going on the market and the GPU actually being shipped to the buyer. Our results are robust to variations in this buffer time—see Section B.4.
@@
The precise way to amortize the training cost through exponential depreciation is:
@@
where training time is in years. However, we could estimate chip-hours more often and more reliably than the training time or the number of chips separately. This is because chip-hours can also be estimated from training compute in FLOP divided by the FLOP/s achieved during training. We used a linear approximation to take advantage of these chip-hour estimates:
Amortized training cost=Start value per chip×Training chip-hours(365×24) $\text{Amortized training cost}=\text{Start value per chip}\times\frac{\text{Training chip-hours}}{(365\times 24)\text{ hours/year}}\times r\ln{10}$
This approximation is valid if (Training $(\text{Training time})\times r\ln{10}$ is small, and this is the case for the training times in our data and our choice of $r=0.14$. In an extreme case, a training time of 1 year results in $1\times 0.14\ln(10)\sim=32\%$ deprecation compared to $1-\exp(-1\times 0.14\ln(10))\sim=28\%$ depreciation. This is not a large difference relative to other sources of uncertainty.
Due to NVIDIA covering defects and component failures under warranty, we concluded that hardware failures are not a significant source of depreciation relative to hardware progress. As one data point, an average of 1 to 2 failures per week occurred when training the BLOOM model on a cluster of 384 NVIDIA A100 GPUs [25]. Even if these were all catastrophic failures, the expected hardware lifetime would be 3.7 years. We expect that NVIDIA replaces or repairs defective GPUs on a faster timescale, which makes the cost of failure small compared to hardware price depreciation.
@@ A.5 Cloud price selection
To estimate training costs from cloud rental prices, we matched the hardware type and publication date of each ML model with a price from the hardware price database. To do this, we first filtered the database for prices which matched the hardware type and the most likely cloud provider that would be used, with the latter based on the developer of the ML model, e.g. using Google Cloud for any Google lab, and using Microsoft Azure for OpenAI based on their partnership with Microsoft. We then estimated the date of hardware procurement as the publication date minus the training time, and minus a further 2 months to account for preparation before the training run. [^5] The choice of 2 months was an educated guess. The matching of models to cloud prices was not very sensitive to this choice because the price data was sparse and stable over time. If the training time was unavailable, we used the median value of approximately 33 days.
@@ A.6 Accounting for compute used throughout model development
One source of evidence on the allocation of compute is the training of smaller model sizes for a given architecture. For example, smaller versions of GPT-3 used 4.5e22 FLOP (based on $compute=6\times parameters\times tokens$[14, Table 2.1]. This shows at least 14% of compute was spent outside the main training run. Similar reasoning for BLOOM reveals about 63% of compute was used on smaller models [25, Table 5].
@@ B.5 Excluding equity from R&D staff costs
To measure the impact of equity on the total amortized model development cost, Figure 9(a) shows the cost breakdown with equity excluded from the R&D staff cost. The proportion of cost on R&D staff decreases from 29–49% with equity included, to 19–33% with equity excluded.
Figure 9: (a) Breakdown of total amortized model development costs for selected models, with equity excluded from the R&D staff cost. Hardware costs are amortized to the total number of chip-hours spent on experiments and training, while R&D staff costs cover the duration of development from initial experiments to publication. Error bars indicate 90% credible intervals, while the main bar values are medians. (b) Costs components as a percentage of the total, based on median estimates.
@@ Appendix C Power capacity for model training
where “Hardware TDP” includes all server hardware. We find a growth rate of 2.2x per year (90% CI: 1.9x to 2.6x). Gemini Ultra has the largest estimated power capacity, at around 35 MW. Projecting the trend forward from Gemini Ultra, the most power-intensive training run would draw 1 GW at some point in 2028. To put this in context, the top ten largest power plants in the United States have a capacity ranging from 3 GW to 7 GW [38].